# GVEL: Fast Graph Loading in Edgelist and Compressed Sparse Row (CSR) formats

### v3

@@ Abstract.
Efficient IO techniques are crucial in high-performance graph processing frameworks like Gunrock and Hornet, as fast graph loading can help minimize processing time and reduce system/cloud usage charges. This research study presents approaches for efficiently reading an Edgelist from a text file and converting it to a Compressed Sparse Row (CSR) representation. On a server with dual 16-core Intel Xeon Gold 6226R processors and MegaRAID SAS-3 storage, our approach, which we term as GVEL, outperforms Hornet, Gunrock, and PIGO by significant margins in CSR reading, exhibiting an average speedup of 78×78$\times$78 ×, 112×112$\times$112 ×, and 1.8×1.8$\times$1.8 ×, respectively. For Edgelist reading, GVEL is 2.6×2.6$\times$2.6 × faster than PIGO on average, and achieves a Edgelist read rate of $1.9$ billion edges/s. For every doubling of threads, GVEL improves performance at an average rate of 1.9×1.9$\times$1.9 × and 1.7×1.7$\times$1.7 × for reading Edgelist and reading CSR respectively.
@@ 2. Related work
Incorporation of Memory-Mapped I/O holds promise for improving graph loading efficiency of frameworks. A number of research studies have worked on optimizing the performance of memory mapped IO. The approaches include using a lightweight userspace memory service (Li et al., 2019), huge pages (Malliotakis et al., 2021), and asynchronous techniques (Imamura and Yoshida, 2019). A large number of works have optimized memory mapped IO for fast low-latency storage devices (Song et al., 2012, 2016; Papagiannis et al., 2020, 2021; Alverti et al., 2022; Leis et al., 2023). Essen et al. (Van Essen et al., 2015) and Feng et al. (Feng et al., 2023) focus on enabling programs to efficiently process out-of-core datasets through memory mapping.
@@ 4.1. Reading Edgelist from text file
Among the approaches given in Sections 4.1.1-4.1.6, we find using mmap() with custom number parsers (mmap-custom) to be the best approach. The pseudocode for mmap-custom is given Algorithm 1. It loads per-thread Edgelists from a file with the best performance, and is enacpsulated in the readEdgelist() function. First, the counts of edges read per thread ($counts$) are initialized, and the components of per-thread Edgelists are obtained ($edges$) in lines 12-13. This is followed by a loop (lines 15-44), where each iteration processes a block of characters (of size $\beta=256K$) in the text file in parallel across different threads. The text file is assumed to have been memory mapped as $data$. Inside the loop, $j$ keeps track of the number of edges processed by the current thread. In line 17, the getBlock() function is called to retrieve the begin and end of current block of text ($[b,B]$), which is processed in the main loop (lines 18-18). The main loop parses edges (lines 20-28), adjusts vertex indices to be zero-based (line 30), and adds them to the Edgelist of the current thread while updating per-partition vertex degrees with atomic operations (lines 32-36). Reverse edges are added for symmetric graphs (lines 38-43). Finally, counts of processed edges per thread are recorded (line 44) and returned (line 45).
The getBlock() function retrieves a block of characters to process, starting from index $i$. It ensures that the block starts and ends on newline characters for proper parsing. The block size is determined by the parameter $\beta$, which is set to $256K$.
@@ 4.2.3. Obtain CSR from global vertex degrees (csr-global)
Now that we have the degrees, we need to combine the per-thread Edgelists into a CSR data structure. In this approach, we first obtain global vertex degrees along with reading per-thread Edgelists (as with degree-global), and then convert the per-thread Edgelists to a global CSR in parallel using atomic operations. As shown in Figure 3, this approach has poor performance (it is 11.4×11.4$\times$11.4 × slower than simply reading Edgelists).
@@ 4.2.4. Obtain CSR from 4-partitioned vertex degrees (csr-partition4)
Among these, we find that obtaining CSR from vertex degrees, where the vertex degrees are computed in four separate partitions using $\bmod$ operator with the current thread id, performs the best. The pseudocode for csr-partition4 is given in Algorithm 2. It transforms an Edgelist representation of a graph into CSR format. First, the components of global CSR ($csr$), per-partition CSRs ($pcsr$), and per-thread Edgelists ($edges$) are obtained in lines 12-14. The algorithm then computes offsets for each partition in parallel (lines 16-16) using exclusive scan operations, and updates the per-partition CSRs concurrently (lines 19-26). Atomic operations ensure thread safety when updating the matrices. In lines 28-31, per-partition offsets are fixed as they were updated during the edge insertion process above. The algorithm then combines per-partition degrees (lines 33-35), computes global offsets (line 37), and merges the per-partition CSRs into the global CSR (lines 39-39).
@@ 5.1. Experimental Setup
We use a server that has two $16$-core x86-based Intel Xeon Gold 6226R processors running at $2.90$ GHz. Each core has an L1 cache of $1$ MB, an L2 cache of $16$ MB, and a shared L3 cache of $22$ MB. The machine has $93.4$ GB of system memory and runs on CentOS Stream 8. We use GCC 8.5 and OpenMP 4.5. Table 1 shows the graphs we use in our experiments. All of them are obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019).
@@ 5.3. Performance Comparison for Reading Edgelist
In Figure 7, we compare the performance of GVEL for reading Edgelist with PIGO. As before, we measure the Edgelist reading time with each framework five times, for averaging. GVEL is on average 2.6×2.6$\times$2.6 × faster than PIGO - with little skew, i.e., GVEL is more or less equivalently faster than PIGO on all graphs. This is more probably than not, due to the reading of Edgelist being pleasingly parallel. On the sk-2005 graph, GVEL achieves a reading performance of $1.9$ billion edges/s.

### v5

@@ Abstract.
Efficient IO techniques are crucial in high-performance graph processing frameworks like Gunrock and Hornet, as fast graph loading can help minimize processing time and reduce system/cloud usage charges. This research study presents approaches for efficiently reading an Edgelist from a text file and converting it to a Compressed Sparse Row (CSR) representation. On a server with dual 16-core Intel Xeon Gold 6226R processors and Seagate Exos 10e2400 HDDs, our approach, which we term as GVEL, outperforms Hornet, Gunrock, and PIGO by significant margins in CSR reading, exhibiting an average speedup of 78×78$\times$78 ×, 112×112$\times$112 ×, and 1.8×1.8$\times$1.8 ×, respectively. For Edgelist reading, GVEL is 2.6×2.6$\times$2.6 × faster than PIGO on average, and achieves a Edgelist read rate of $1.9$ billion edges/s. For every doubling of threads, GVEL improves performance at an average rate of 1.9×1.9$\times$1.9 × and 1.7×1.7$\times$1.7 × for reading Edgelist and reading CSR respectively.
@@ 2. Related work
Incorporating Memory-Mapped I/O holds promise for improving graph loading efficiency of frameworks. A number of research studies have worked on optimizing the performance of memory mapped IO. The approaches include using a lightweight userspace memory service (Li et al., 2019), huge pages (Malliotakis et al., 2021), and asynchronous techniques (Imamura and Yoshida, 2019). A number of works have optimized memory mapped IO for fast low-latency storage devices (Song et al., 2012, 2016; Papagiannis et al., 2020, 2021; Alverti et al., 2022; Leis et al., 2023). Essen et al. (Van Essen et al., 2015) and Feng et al. (Feng et al., 2023) focus on enabling programs to efficiently process out-of-core datasets through memory mapping.
@@ 4.1. Reading Edgelist from text file
Among the approaches given in Sections 4.1.1-4.1.6, we find using mmap() with custom number parsers (mmap-custom) to be the best approach. The pseudocode for mmap-custom is given Algorithm 1. It loads per-thread Edgelists from a file with the best performance, and is enacpsulated in the readEdgelist() function (lines 11-45). First, the counts of edges read per thread ($counts$) are initialized, and the components of per-thread Edgelists are obtained ($edges$) in lines 12-13. This is followed by a loop (lines 15-44), where each iteration processes a block of characters (of size $\beta=256K$) in the text file in parallel across different threads. The text file is assumed to have been memory mapped as $data$. Inside the loop, $j$ keeps track of the number of edges processed by the current thread. In line 17, the getBlock() function is called to retrieve the begin and end of current block of text ($[b,B]$), which is processed in the main loop (lines 18-43). The main loop parses edges (lines 20-28), adjusts vertex indices to be zero-based (line 30), and adds them to the Edgelist of the current thread while updating per-partition vertex degrees with atomic operations (lines 32-36). Reverse edges are added for symmetric graphs (lines 38-43). Finally, counts of processed edges per thread are recorded (line 44) and returned (line 45).
The getBlock() function (lines 47-54) retrieves a block of characters to process from the memory-mapped file, starting from index $i$. It ensures that the block starts and ends on newline characters for proper parsing. The block size is determined by the parameter $\beta$, which is set to $256K$.
@@ 4.2.3. Obtain CSR from global vertex degrees (csr-global)
Now that we have the degrees, we need to combine the per-thread Edgelists into a CSR. In this approach, we first obtain global vertex degrees along with reading per-thread Edgelists (as with degree-global), and then convert the per-thread Edgelists to a global CSR in parallel using atomic operations. As shown in Figure 3, this has poor performance (it is 11.4×11.4$\times$11.4 × slower than simply reading Edgelists).
@@ 4.2.4. Obtain CSR from 4-partitioned vertex degrees (csr-partition4)
Among these, we find that obtaining CSR from vertex degrees, where the vertex degrees are computed in four separate partitions using $\bmod$ operator with the current thread id, performs the best. The pseudocode for csr-partition4 is given in Algorithm 2. It transforms an Edgelist representation of a graph into CSR format. First, the components of global CSR ($csr$), per-partition CSRs ($pcsr$), and per-thread Edgelists ($edges$) are obtained in lines 12-14. The algorithm then computes offsets for each partition in parallel (lines 16-17) using exclusive scan operations, and updates the per-partition CSRs concurrently (lines 19-26). Atomic operations ensure thread safety when updating the matrices. In lines 28-31, per-partition offsets are fixed as they were updated during the edge insertion process above. The algorithm then combines per-partition degrees (lines 33-35), computes global offsets (line 37), and merges the per-partition CSRs into the global CSR (lines 39-48).
@@ 5.1. Experimental Setup
We use a server that has two $16$-core x86-based Intel Xeon Gold 6226R processors running at $2.90$ GHz. Each core has an L1 cache of $1$ MB, an L2 cache of $16$ MB, and a shared L3 cache of $22$ MB. $12$ Seagate Exos 10e2400 $2.4$ TB SAS-12Gbps 2.5inch HDDs (DL2400MM0159), each with a rotational speed of $10k$ rpm, a sustained transfer rate of $241$ (outer diameter) to $117$ Mbps (inner diameter), and an average latency of $2.9$ ms, are installed on the server — connected with a MegaRAID SAS-3 3008 SAS+SATA Controller. The machine has $512$ GB of system memory, and runs on CentOS Stream 8. We use GCC 8.5 and OpenMP 4.5. Table 1 shows the graphs we use in our experiments. All of them are obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019).
@@ 5.3. Performance Comparison for Reading Edgelist
In Figure 7, we compare the performance of GVEL for reading Edgelist with PIGO. As before, we measure the Edgelist reading time with each framework five times, for averaging. GVEL is on average 2.6×2.6$\times$2.6 × faster than PIGO - with little skew, i.e., GVEL is more or less equivalently faster than PIGO on all graphs. This is more probably than not, due to the reading of Edgelist being pleasingly parallel. On the sk-2005 graph, GVEL achieves a read performance of $1.9$ billion edges/s. Figure 8 illustrates the time GVEL requires to read the edge-list and convert it to CSR for each graph in the dataset, presented separately.
@@
Figure 8. Time taken by GVEL for edge-list and CSR loading on 13 different graphs.